OpenStudy (anonymous):

A boy standing on the ground 25 feet from the base of a building looks up at the top of a building. The angle of elevation is 72°. Disregarding the boy’s height, how tall is the building? Round your answer to the nearest tenth of a foot.

OpenStudy (anonymous):

7.55590

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

\[25\tan(72)\] whatever that is. it is not 7.5 though. not a reasonable height for a building

OpenStudy (anonymous):

\[25 \text{Tan}[\pi 72/180]=25 \sqrt{5+2 \sqrt{5}}=76.9 \]

OpenStudy (anonymous):

that is one way to do it. another is to whip out a calculator, make sure it is in degree mode, and compute

OpenStudy (anonymous):

I thank you your comment. Have been converting degrees to radians in Mathematica until now. Was unaware that there was a degree version of the argument. \[\text{Tan}[72{}^{\circ}] \]computes just like the radian version of the argument. The handheld calculator is probably the fastest in any case; three key strokes, 7,2,Tan. Just bought a fresh set of batteries for my HP 50g. I may have missed out on some medals in the past.